OpenStudy (moongazer):

(2/3+5/7)/(2/3-5/7) how to solve this the easiest way?

OpenStudy (saifoo.khan):

solve the brackets!

OpenStudy (saifoo.khan):

them multiply them!

OpenStudy (moongazer):

my result is always -1 isn't it not correct

OpenStudy (moongazer):

please solve it for me

OpenStudy (zarkon):

\[-\frac{29}{441}\]

OpenStudy (anonymous):

-0.06

OpenStudy (anonymous):

\[(2/3+5/7)/(2/3-5/7) \] \[=((14+15)/21) /((14-15)/21)\] \[=29/-1\] \[=-29\]

OpenStudy (zarkon):

\[\frac{14+15}{21}\frac{14-15}{21}=\frac{29(-1)}{21\cdot 21}=\frac{-29}{441}\]

OpenStudy (zarkon):

oops...thought they were doing multiplication